Consider the following information for Stocks A, B, and C. The returns on the three stocks are positively correlated, but they are not perfectly correlated.

Stock Expected Return Standard Deviation Beta
A 9.55% 15% 0.9
B 10.45% 15% 1.1
C 12.70% 15% 1.6

Fund P has one-third of its funds invested in each of the three stocks. The risk-free rate is 5.5%, and the market is in equilibrium.

a. What is the market risk premium?

b. What is the beta of Fund P?

c. What is the required return of Fund P?

d. Would you expect the standard deviation of Fund P to be less than 15%, equal, or greater than?
Answer #1

a.

Expected Return = (0.0955 + 0.1045 + 0.1270)/3 = 10.90%

Beta = (0.90 + 1.10 + 1.60)/3 = 1.2

Using CAPM model,

Market Risk Premium = 0.1090 - 0.055 = 5.40%

b.

Beta = 1.2

c.

Required Return = 10.90%
Answer #2
(a) Re= Rrf+(Rm-Rrf)B 9%= 5.5%+(Rm-5.5%)0.8 (Rm-5.5%)= 9%-5.5%/0.8 Rm= 0.04375+5.5% Rm= 9.875% (b) B(Q)= 0.333(0.8)+0.333(1.2)+0.333(1.6) B(Q)= 1.1964 (c) Required return of beta fund (Q) = 5.5%+(9.875%-5.5%)1.2 =5.5%+0.0525 =10.75%
source: Teacher
answered by: anonymous